I need to find out the total amount of sales. My sales tax of 6% = $2,357.10.
What formula is used?
Thank you.

Answer by tutorcecilia(2144)

Use the formula for percentages:
.
Part/Whole = % / 100 [Plug-in the values]
(2357.10)/x = 6%/100 [Cross-multiply and solve for s]
(2357.10)(100) = (x)(6%)
235710 = 6x
235710/6 = 6x/6
39285 = x
.
Plug (x=39285) back into the original equation to verify the answer:
Part/Whole = % / 100
(2357.10)/(39285) = 6%/100
.06 = .06 [Checks out]